OpenStudy (anonymous):

The length of a flat screen TV is x2 + 1 feet. The width is x + 3 feet. What algebraic expression can be used to represent the area of the TV?

OpenStudy (anonymous):

Area = Length x Width so \[A = (x^2+1)(x+3) = x^3 + 3x^2 + x + 3.\]

OpenStudy (anonymous):

thanks!